OpenStudy (anonymous):

solve by the substitution method 8x+7y=-6 -8x+y=54

jhonyy9 (jhonyy9):

8x+7y=-6 -8x+y=54 --- y=54+8x 8x+7(54+8x)=-6 8x+378+56x=-6 64x=-384 /4 16x=-96 /4 4x=-24 /4 x=-6 --- y=54+8(-6)=54-48=6 x=-6 y=6